The ball circumference was measured and the result was p = 38.77 ± 0.20cm. By what accuracy can the ball surface be calculated?
1
Expert's answer
2019-12-16T11:34:39-0500

The ball circumference is: "p = 38.77 \space \text{cm}"

The error of measurements is: "\Delta p = 0.2 \space \text{cm}"

The surface can be calculated by the formula: "S = \frac{1}{4\pi}p^2"

Taking the differential and keeping only linear part, find the error of surface calculations:

"\Delta S = \frac{1}{2\pi}p\cdot\Delta p"

Pluging numerical values obtain :

"\Delta S = 1.2 \space \text{cm}^2"

So, the surface can be measured with accuracy: "\pm1.2 \space \text{cm}^2"